Butane, C4H10, is an easily liquefied gaseous fuel. Calculate the density of butane gas at 0.897 atm and 24°C. Give the answer in grams per liter


1
Expert's answer
2021-05-31T02:05:26-0400

"T = 273+24=297 K."


"R= 0.082"


"P=0.897" atm


PV = RT


"V=\frac{RT}{P}"


"V=\frac{0.082×297}{0.897}= 26.97= 26970cm^3/mole" =


Molar mass of "C_4H_{10}= 56g\/mole"


Density "= \frac{56}{26970}= 0.002 g/cm^3"